# MambaTrack: Exploiting Dual-Enhancement for Night UAV Tracking

### v1

@@ I Introduction
Visual tracking, a fundamental task in computer vision, has received widespread attention in various unmanned aerial vehicle (UAV) applications, e.g., aerial photography, target following, smart agriculture, and delivery. It aims to sequentially track a moving object in a video from its initial location. Despite significant advancements, recent efforts have largely concentrated on daytime UAV tracking [1, 2, 3]. Consequently, numerous UAV-related applications are severely hindered due to the insufficient exploration of night UAV tracking [4].
Generally, video frames captured at night have lower brightness and contrast compared to daytime [2, 5, 6]. To pursue better tracking performance in the dark night, current works mainly employ two strategies, i.e., low-light enhancement [7, 8, 9] and unsupervised domain adaptation (UDA) [4, 10]. Among them, ADTrack [9] is a pioneering anti-darkness UAV tracker that adopts a tone mapping algorithm within a correlation filter-based framework to enhance tracking robustness in low-light conditions. Later, some works proposed to eliminate the influence of poor illumination by multiple iterations [7] or a spatial-channel Transformer [8]. However, these low-light enhancement-based methods either struggle to preserve important local details of the image or have a high computational burden. In contrast, UDA-based methods [4, 10] propose to learn from unlabeled data in the target domain (i.e., nighttime scenarios), and utilize knowledge from the source domain (i.e., daytime scenarios) to improve performance in the target domain. However, these methods are challenged by the imbalanced source/target distribution [10] and limited data [11] for night UAV tracking.
@@ III-A Implementation Details
Our experimental platform is a server with 8 RTX A6000 GPUs. We adopt Vim-S [28], GPT-NeoX [29], and Mamba-130M [14] as visual mamba encoder, text tokenizer and language mamba encoder, respectively. The template and search region are $2^{2}$ and $4^{2}$ times of the target and then resized to $128\times 128$ and $256\times 256$. The balance factors $\lambda_{1}$, $\lambda_{GIoU}$, and $\lambda_{focal}$ are set to 5, 2, and 1.5. The visual and language dimensions are 384 and 768. The low-light enhancer is trained using LOL [30]. The tracker is trained using the training splits of TrackingNet [31], GOT-10k [32], LaSOT [33], COCO [34], and WebUAV-3M [1]. We evaluate different trackers on five night UAV tracking datasets, i.e., DarkTrack2021 [8], NAT2021 [4], NAT2021L [4], UAVDark70 [9], and UAVDark135 [2]. For fair comparisons, we primarily adopt tracking results of methods from original datasets [8, 4, 9, 2]. Due to the absence of language annotations in these night UAV tracking datasets, we manually annotate 518 language prompts (each video with a sentence describing the class name of the target, its attributes, and the surrounding environment) for multimodal night UAV tracking. To facilitate future research, we will release these language annotations.
@@ III-D Speed and Model Size Analysis
Inference speed and model parameters are two key considerations since UAV platforms usually require low latency and low power consumption [1, 41]. As shown in Tab. III, MambaTrack achieves a real-time inference speed of 42 frame per second (FPS). Compared with current SOTA CNN-Transformer-based trackers (i.e., STARK50 [35] and VLTTT [37]), and Transformer-based trackers (i.e., JointNLT [12] and CiteTracker [13]), our MambaTrack reduces learnable parameters by 43.6% to 91.0%. Compared with current SOTA VL-based method CiteTracker, our GPU memory usage is reduced by 50.2%.
@@ IV Conclusion
This work introduces MambaTrack, a novel tracker leveraging dual enhancement techniques for night UAV tracking. MambaTrack utilizes a mamba-based low-light enhancer and a cross-modal mamba network to improve tracking performance in low-light conditions. Another significant contribution of this work is that we annotate language prompts for existing datasets, constructing a new vision-language night UAV tracking task. Extensive experiments on five benchmarks demonstrate the superiority and efficiency of the proposed approach compared to current SOTA methods. In the future, we plan to explore mamba-based trackers for more modalities, e.g., depth and thermal infrared.

### v2

@@ I Introduction
Visual tracking, a fundamental task in computer vision, has received widespread attention in various unmanned aerial vehicle (UAV) applications, e.g., aerial photography, target following, smart agriculture, and delivery. It aims to sequentially track a moving object in a video from its initial location. Despite significant advancements, recent efforts have largely concentrated on daytime UAV tracking [1, 2, 3]. Consequently, numerous UAV-related applications are severely hindered due to the insufficient exploration of night UAV tracking [4, 5].
Generally, video frames captured at night have lower brightness and contrast compared to daytime [2, 6, 7, 8]. To pursue better tracking performance in the dark night, current works mainly employ two strategies, i.e., low-light enhancement [9, 10, 11] and unsupervised domain adaptation (UDA) [4, 12]. Among them, ADTrack [11] is a pioneering anti-darkness UAV tracker that adopts a tone mapping algorithm within a correlation filter-based framework to enhance tracking robustness in low-light conditions. Later, some works proposed to eliminate the influence of poor illumination by multiple iterations [9] or a spatial-channel Transformer [10]. However, these low-light enhancement-based methods either struggle to preserve important local details of the image or have a high computational burden. In contrast, UDA-based methods [4, 12] propose to learn from unlabeled data in the target domain (i.e., nighttime scenarios), and utilize knowledge from the source domain (i.e., daytime scenarios) to improve performance in the target domain. However, these methods are challenged by the imbalanced source/target distribution [12] and limited data [13] for night UAV tracking.
@@ III-A Implementation Details
Our experimental platform is a server with 8 RTX A6000 GPUs. We adopt Vim-S [30], GPT-NeoX [31], and Mamba-130M [16] as visual encoder, text tokenizer and language encoder, respectively. The template and search region are $2^{2}$ and $4^{2}$ times of the target and then resized to $128\times 128$ and $256\times 256$. The balance factors $\lambda_{1}$, $\lambda_{GIoU}$, and $\lambda_{focal}$ are set to 5, 2, and 1.5. The visual and language dimensions are 384 and 768. The low-light enhancer is trained using LOL [32]. The tracker is trained using the training splits of TrackingNet [33], GOT-10k [34], LaSOT [35], COCO [36], and WebUAV-3M [1]. We evaluate different trackers on five night UAV tracking datasets, i.e., DarkTrack2021 [10], NAT2021 [4], NAT2021L [4], UAVDark70 [11], and UAVDark135 [2]. For fair comparisons, we primarily adopt tracking results of methods from original datasets [10, 4, 11, 2]. Due to the absence of language annotations in these night UAV tracking datasets, we manually annotate 518 language prompts (each video with a sentence describing the class name of the target, its attributes, and the surrounding environment) for multimodal night UAV tracking. To facilitate future research, we will release these language annotations.
@@ III-D Speed and Model Size Analysis
Inference speed and model parameters are two key considerations since UAV platforms usually require low latency and low power consumption [1, 43]. As shown in Tab. III, MambaTrack achieves a real-time inference speed of 42 frames per second (FPS). Compared with current SOTA CNN-Transformer-based trackers (i.e., STARK50 [37] and VLTTT [39]), and Transformer-based trackers (i.e., JointNLT [14] and CiteTracker [15]), our MambaTrack reduces learnable parameters by 43.6% to 91.0%. Compared with the current SOTA VL-based method CiteTracker, our GPU memory usage is reduced by 50.2%.
@@ IV Conclusion
This work introduces MambaTrack, a novel tracker leveraging dual enhancement techniques for night UAV tracking. MambaTrack utilizes a mamba-based low-light enhancer and a cross-modal mamba network to improve tracking performance in low-light conditions. Another significant contribution of this work is that we annotate language prompts for existing datasets, constructing a new vision-language night UAV tracking task. Extensive experiments on five benchmarks demonstrate the superiority and efficiency of the proposed approach compared to current SOTA methods.
Acknowledgements. This work was supported by the National Natural Science Foundation of China (No. 62471420 and 62101351), and the Major Project of Technology Innovation and Application Development of Chongqing (CSTB2023TIAD-STX0015).